OpenStudy (anonymous):

Solve the system by substitution. 2x + y=-11 3x - 4y=11 i dont get this at all can somone just break this down????

OpenStudy (anonymous):

1st equation : subtract 2x in both sides: you will get y=-11-2x Plug y in 2nd equation: 3x -4(-11-2x) =11 3x+44+8x=11 Solve for x then you can find for y